OpenStudy (anonymous):

A falling stone is at a certain instant 302 feet above the ground. 3 seconds later it is only 14 feet above the ground. From what height was it dropped

OpenStudy (anonymous):

delta h = 14-302 = -Vi*(3) -1/2*g*(3)^2 find Vi use that to find the height it's fallen up to that point by Vi^2 = 2*g*d d+302 = initial height